Solve using the substitution method 0.3x + 0.2y = 5, 0.5x + 0.4y = 11
Answer by jorel1380(3719)

.3x+.2y=5
3x+2y=50
6x+4y=100
4y=100-6x;
.5x+.4y=11
5x+4y=110
4y=110-5x
110-5x=100-6x
x=-10
2y=80
y=40..